Find the numerical measures of the angles of the triangle whose angle measures are given:
x, x + 10°, x + 20°

Answer by laoman(51)

The angle of a triangle sum up to 180 degrees
therefore, x+x+10+x+20 = 180
3*x+30 = 180
3*x = 150
x = 150/3
x=50
therefore the angles are 50 deg, 60 deg and 70 deg
QED